You purchased a house for $850000 cash 4 years ago. You can sell it today for $980000. What rate of return did you earn on this investment? Round your answer to the nearest tenth of a percent.              

Options:

115.3%                                                       

none of the choices

15.3%

3.1%           
Answer #1

RATE OF RETURN ON INVESTMENT = [(SELLING PRICE/PURCHASE PRICE) - 1] X 100  

RATE OF RETURN ON INVESTMENT = [(980000/850000) - 1] X 100 = (1.1529- 1) X 100 = 15.3%

NOTE :THEY HAVE NOT ASKED ANNUAL RATE OF RETURN.

ANSWER : 15.3% (Thumbs up please)